A straight line is parallel to y=-3x+2 and passes through (2,-1).
1)Find the gradient of the straight line
2) The equation of the straight line
3) The y-intercept of the straight line

Answer by richwmiller(17219)

1) it will be the same as in the equation y=-3x+2. the gradient or slope is -3
2) y=-3x+b
plug in x and y to find b
-1=-3*2+b
-1=-6+b
add 6
5=b which is the y intercept
y=-3x+5